Hockey Teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 69 points. They won 12 more games than they tied. How many wins and how many ties did the team have?
How many wins __
How many ties __
I'm sorry, but i am just drawing a total blank today. Thank you all.

Answer by stanbon(75887)

Hockey Teams receive 2 points when they win and 1 point when they tie.
One season, a team won a championship with 69 points.
They won 12 more games than they tied.
How many wins and how many ties did the team have?
------------------------------------
Equations:
Quantity Equation: w = t + 12
Points Equation..: 2w + 1t = 69
-----------------------------------
Substitute to solve for "t":
2(t+12) + t = 69
3t + 24 = 69
3t = 45
t = 15 (# of ties)
Since w = t+12,w = 15+12 = 27 (# of wins)
=====================================
Cheers,
Stan H.